Plese help me understand this problem, I know that there should be $10.75 left after the purchase but how would i figure out the rest of the problem?
Problem:
A student makes a $9.75 purchase at the bookstore with a $20 bill. the store has no bills and gives the change in quarters and fifty cent pieces. their are 30 coins in all. how many of each kind are there? How many quarters are there in the change?

Found 2 solutions by Alan3354, stanbon:
Answer by Alan3354(69443)

A student makes a $9.75 purchase at the bookstore with a $20 bill. the store has no bills and gives the change in quarters and fifty cent pieces. their are 30 coins in all. how many of each kind are there? How many quarters are there in the change?
---------------
The change is $10.25.
q = quarters
h = half-dollars
q + h = 30 (30 coins) --> q = 30-h
25q + 50h = 1025
Sub for q
25(30-h) + 50h = 1025
750-25h + 50h = 1025
25h = 275
h = 11
q = 19

Answer by stanbon(75887)

A student makes a $9.75 purchase at the bookstore with a $20 bill. the store has no bills and gives the change in quarters and fifty cent pieces. their are 30 coins in all. how many of each kind are there? How many quarters are there in the change?
---------------
The change = 20-9.75 = 10.25
------------------
Equations:
Quantity: Q + F = 30
Value::::25Q+50F = 1025 cents
--------------------------------------
Modify the 1st equation to prepare for elimination:
25Q + 25F = 25*30
---
Subtract that from 25Q + 50F = 1025 to solve for "F":
25F = 1025-25*30
25F = 275
F = 11 (# of 50 cent pieces)
--------------
Substitute into Q + F = 30 to solve for Q:
Q = 19 (# of quarters)
========================================
Cheers,
Stan H.